# QOS: A Quantum Operating System

### v1

@@ 3.1. Fidelity
Figure 4. (a) Challenge #2, Temporal variance (§ 3.2) Fidelity of a 6-qubit GHZ circuit on IBM Perth, across 120 calibration days. There are 20 pairs of days with more than 5% difference in fidelity. (b) Challenge #3, Utilization (§ 3.3) Maximum utilization achieved on a 27-qubit QPU for nine benchmarks while maintaining at least 0.75 fidelity. The average utilization is $26.3\%$, and the max is $29.6\%$. (c) Challenge #4, QPU Load (§ 3.4) Number of pending jobs on different IBM QPUs. The groups separated by vertical red lines indicate QPUs of the same size. There is up to 57$\times$ difference in number of jobs between QPUs of the same size.
Results. Our results are highlighted in Figure 3 (a). The x axis depicts the circuit size as the number of qubits while the y axis shows the fidelity, where higher is better. The experiment is run on the IBM Kolkata 27-qubit QPU. For each increase in qubits, the average fidelity decreases, up to $98.9\%$ from 4 to 24 qubits. Moreover, it is physically impossible to run circuits with a size larger than 27 qubits, since we cannot map them.
Implication. NISQ devices are limited due to size and noise and, therefore, cannot be practically used for large quantum circuits; either logically, because the circuit doesn’t fit in the device, or the execution results would be convoluted from noise errors, which translates to low fidelity.
Key Idea #1: Circuit Optimizations: To increase fidelity, we need a generic optimization infrastructure that transforms circuits into a physically and practically executable size.
3.2. Spatial and Temporal Heterogeneity
In the classical domain, two identical CPUs perform similarly for all applications, and at each point in time. In contrast, QPUs exhibit differences in the layout and connectivity of qubits (Gyongyosi and Imre, 2019) and variations in noise errors even for QPUs of the same model, which leads to spatial performance variance. Moreover, QPUs are calibrated regularly (§ 2.2), and after each calibration, the noise properties change (Wittler et al., 2021). As a result, the execution fidelity can vary across different calibration cycles, leading to temporal performance variance.
Results. Figure 3 (b) shows a 12-qubit GHZ circuit’s fidelity on different IBM QPUs. Fidelity varies across the QPUs, with a maximum difference of 38% from best to worst. Note that all six QPUs are of the same model (Falcon r5.11).
Figure 4 (a) shows a 6-qubit GHZ circuit’s fidelity over 120 calibration days executed on the IBM Perth 7-qubit QPU, where each data point represents a single day’s fidelity. The largest single-day difference in fidelity is 96.5%, and there are 20 instances of a single-day fidelity drop of more than 5%. Note that there is no way of predicting a QPU’s future calibration data to expect such performance differences.
@@ 3.2. Spatial and Temporal Heterogeneity
Key Idea #2: Performance Estimation: We estimate a circuit’s potential performance on the available QPUs to automatically select the best-performing candidate(s).
3.3. Utilization
@@
Results. Figure 4 (b) shows the maximum utilization of the IBM Kolkata 27-qubit QPU for nine benchmarks while maintaining at least 0.75 fidelity. No benchmark exceeds 30% utilization, while the average is 26.3%. Higher fidelity values would yield even lower utilization and vice-versa.
Implications. There is a tradeoff between QPU utilization and performance (fidelity). In general, the lower utilization, the higher fidelity, and vice-versa. In contrast to the classical domain, the tension between these objectives is vastly larger.
Key Idea #3: Multi-programming: We spatially multiplex quantum circuits to increase system utilization (also known as multi-programming (Das et al., 2019)), and when combined with circuit optimizations, it also increases fidelity.
3.4. QPU Load Imbalance
The quantum cloud faces QPU load imbalance. The root cause is spatiotemporal heterogeneity (§ 3.2), combined with the manual QPU selection offered by the current quantum cloud model (ibmQuantum, [n. d.]). This leads to users selecting the “best performant” QPU based on empirical or arbitrary metrics (Ravi et al., 2021a).
@@ 3.4. QPU Load Imbalance
Implications. Load imbalance leads to long waiting times for the users and thus, low quality of service. Additionally, there is no 1-1 mapping between the load and performance differences between QPUs. For instance, the 12-qubit GHZ circuit in Figure 3 (b) performs 1.1$\times$ better on IBM Hanoi than IBM Cairo, yet the former exhibits 57$\times$ higher load.
Key Idea #4: Load-aware Scheduling: We schedule (temporally multiplex) quantum circuits in a load-aware manner to balance the tradeoff between fidelity and waiting times.
4. Overview
Quantum computing is characterized by four main challenges that limit its practicality: (1) Execution fidelity is hindered by the small and noisy QPUs. (2) In contrast to classical accelerators, QPUs exhibit vast spatiotemporal heterogeneities, which renders their performance non-deterministic in both dimensions. (3) QPUs are heavily underutilized to give high-fidelity results. (4) QPUs face vast load imbalance, which leads to prolonged waiting times for the users.
Existing work is narrow-scoped and focuses on tackling one challenge at a time, but unfortunately, there are two main issues with this point solution approach. Firstly, composing the individual mechanisms to address all challenges at once is impossible without a common and unified infrastructure. Secondly, without synergies between the individual mechanisms, it is hard to maximize the objectives of the users, i.e., high fidelity and low waiting times, and the objectives of the quantum cloud operator, i.e., resource efficiency.
To this end, we propose QOS, an end-to-end system that tackles the challenges of quantum computing holistically. QOS strives for three design goals: (1) A unified architecture that supports compiler and OS mechanisms with pluggable policies and tunable configuration for managing the tradeoffs of QC. (2) QOS should enable the execution of large quantum circuits with high fidelity and scale with increasing incoming workloads and additional QPUs. (3) QOS should be resource efficient by achieving high QPU utilization and balancing QPU load to minimize waiting times.
Figure 5. QOS overview (§ 4): QOS consists of two main components: the QOS compiler (top) and the QOS runtime (bottom). Below QOS lie the QPU devices and classical nodes.
4.1. The QOS Architecture
@@
Figure 5 shows the overview of our system’s design. QOS comprises a layered architecture that consists of two main components: the QOS compiler (top) and the QOS runtime (bottom), which we detail next.
@@
Qernel Abstraction. QOS implements a wide range of mechanisms with different abstraction requirements, from the compilation to the execution runtime level. To enable the composability of these mechanisms in a unified architecture, we propose the Qernel abstraction that acts as a common denominator for the QOS mechanisms to apply their policies.
QOS Compiler. We propose the QOS compiler (Figure 5, top), a modular, extensible, and composable compiler infrastructure. It comprises three stages: (1) The frontend of the compiler, the analyzer (§ 5.3), accepts quantum circuits and lifts them to the Qernel abstraction, generates the intermediate representation (IR), and performs IR analysis passes to generate the IR static properties required by the next stages. (2) The middle-end, the optimizer (§ 6), is an extensible and composable set of optimization passes that leverages the IR and static properties to improve the execution fidelity of the quantum circuits with manageable overheads. (3) The backend, the virtualizer (§ 7), compiles the optimized Qernels for the target QPUs, similar to classical target code generation.
QOS Runtime. We propose the QOS runtime (Figure 5, bottom), a system that abstracts away the underlying heterogeneity and balances the tradeoff between the conflicting objectives of the cloud operator (resource efficiency) and the users (high fidelity and low waiting times). The runtime comprises four components: (1) The estimator predicts the fidelity of executing the optimized Qernels to guide scheduling decisions. (2) The multi-programmer, given the estimations, bundles low utilization Qernels to increase QPU utilization. (3) The scheduler multiplexes and runs the Qernels across space and time with the objective to maximize fidelity and minimize waiting times. Finally, (4) the knitter post-processes the Qernel execution results to return the final result to the user.
Figure 6. Compiler Frontend: Analyzer (§ 5). (a) An example quantum circuit of 4 qubits and 5 gates. (b) The Qernel intermediate representation (QIR) (§ 5.1). (c) The refined QIR (§5.1). (d) The Qernel’s static and dynamic properties (§5.2). During compilation, the dynamic properties are void, but they are initialized and used during the runtime.
4.2. Execution Workflow
@@
First, users submit a circuit along with their optimization target and budget 1. The former represents the desired post-compilation circuit size, and the latter quantifies the additional overheads the user is willing to pay. The compiler’s frontend lifts the circuit to the Qernel abstraction and generates the IR and its static properties 2. The middle-end optimizes the Qernels through a modular set of passes 3, then the backend generates the target QPU-optimized Qernels and submits them to the QOS runtime 4. The estimator predicts the fidelity of running the Qernel(s) on the QPUs to guide scheduling 5. The multi-programmer bundles Qernels with low utilization and sends them to the scheduler 6. The scheduler assigns and runs the bundled Qernels, optimizing for maximal fidelity and minimal waiting times 7. After the execution, the bundled results are retrieved by the multi-programmer 8 to be unbundled into separate results and are sent to the knitter 9. Finally, the knitter post-processes the separated results pass and returns them to the user 10.
5. Compiler Frontend: Qernel & Analyzer
5.1. The Qernel Abstraction
The Qernel is the unified abstraction acting as a common denominator for the QOS components. Specifically, a Qernel contains (1) the graph-based IR used by the QOS compiler and (2) the Qernel properties, which comprises static IR properties and dynamic properties used by the QOS runtime.
Qernel Intermediate Representation (QIR). Existing optimization techniques operate at the gate level, analogous to the instruction level in the classical domain. Therefore, we propose a graph-based Qernel Intermediate Representation (QIR) that captures the control flow of a quantum program, similar to the control flow graph of classical programs. By traversing the QIR, the compiler can identify important optimization opportunities, such as pairs of gates that cancel each other (like dead code elimination), gate dependencies (useful for gate scheduling, similar to instruction scheduling), or opportunities to remove hotspot gates (gates that contribute to noise errors in the computation). An example QIR is shown in Figure 6 (b), where the quantum circuit consisting of four qubits and five gates is lifted to the QIR.
Formally, a QIR is a directed acyclic graph (DAG) $G=(V,E)$, where $V$ is the set of gates and every edge $e_{i}\in E$ is the qubit the gate acts on. The edges’ directions reflect dependencies $D=\{(V_{i},V_{j})\in V\times V\}$ between gates, i.e., $V_{i}$ must be scheduled before $V_{j}$. To identify hotspot nodes, we compute the degree of a node $deg(V_{i})$, which reflects the number of control flow paths the gate $V_{i}$ is part of. In the example of Figure 6, the QIR reveals four layers of gates: $l_{1}:(g_{0},g_{1})$, $l_{2}:g_{2}$, $l_{3}:(g_{3},g_{4})$, and $l_{4}:M$, which means that they have to be scheduled in this order, and the pairs in the same layers are susceptible to crosstalk noise (§ 2). Finally, the gate $g_{2}$ is a hotspot node since its degree is 4, the highest in $V$ (the measurements $M$ are terminal nodes and do not count).
Refined QIR. Various optimizations require a simplified representation that captures only the connectivity between qubits. For instance, the connectivity structure might reveal hotspot qubits (Ayanzadeh et al., 2023) that can be removed for fidelity improvement or opportunities for circuit cutting (§ 6). Figure 6 (c) shows the refined QIR, where we can see that $q_{1}$ and $q_{2}$ are only connected by a single gate, and removing it would split the circuit into two smaller circuits. Formally, a refined QIR is an acyclic, undirected, and weighted graph $G=(V,E)$, where $V$ is the set of qubits and every edge $e_{i}\in E$ between two qubits $V_{j},V_{k}$ has a weight $w_{i}\in\mathbb{N}$ that represents the number of gates that act on $V_{j}$ and $V_{k}$.
5.2. Qernel Properties
For applying the diverse set of its mechanisms, QOS requires data structures that keep up-to-date information about the quantum programs. Such information includes (1) the static properties, which are useful for the compiler, and (2) dynamic properties, which are useful for the runtime.
Static Properties. Apart from the IR, optimization passes leverage circuit properties to be more efficient and effective. The properties include the circuit’s size (number of qubits), depth, non-local gates and their types, the number of measurements, and others (Figure 6 (d)). Additionally, we include the features vectors defined in (Tomesh et al., 2022) since they are potentially useful for heuristic-based optimizations or regression-based prediction models (Quetschlich et al., 2023).
Dynamic Properties. The Qernel also contains dynamic properties required by the QOS runtime. These include the Qernel’s execution status (done, failed, running, scheduled), the estimator’s output, i.e., fidelity estimations (§ 8.1), and the final post-processed results (Figure 6 (d)).
@@ 5.2. Qernel Properties
Figure 7. Optimization workflow (§ 6). The initial refined QIR has 7 qubits and 12 gates. (a) The HotSpotNodePass identifies $q_{3}$ as a hotspot node with a degree of 6. (b) We optimize the refined QIR with $s=2,b=3$ by applying the QubitFreezingPass which reduces the qubit and gate counts to 6. (c) We apply the GateCuttingPass to further remove 2 gates. This gives two fragments of 3 qubits each. (d) We have depleted the budget, so we use the QubitReusePass to achieve the goal of $s=2$.
5.3. Frontend: Analyzer
To discover and leverage optimization opportunities, we first need to perform circuit analysis, similar to classical program analysis. The analyzer transforms a quantum circuit into a Qernel and comprises an extensible set of passes that generate the QIR and static properties of the Qernel, which are then used by the optimizer and subsequently by the runtime.
QIR Transformation Pass. The first step in program analysis and optimization is to generate the QIR, implemented by the QIR transformation pass. Figure 6 (a)-(b) shows the generation of the QIR for an example quantum circuit. To generate the QIR, the pass iterates over each logical qubit of the circuit and each gate acting on that qubit. For each such gate, it creates a QIR vertex $g_{i}$ and sets the qubits $q_{j},q_{k}$ the gate acts on as the edges $e_{j},e_{k}$ of $g_{i}$. By convention, the direction of the edge follows the direction from the control to the target qubits. When reaching measurement operations, it simply adds the terminal nodes $M$. This process is repeated until all circuit qubits and gates are covered.
QIR Refinement Pass. To generate the refined QIR (Figure 6 (c)), we implement a transformation pass that traverses the QIR in a depth-first manner (breadth-first is equivalent). For each QIR node visited, i.e., a vertex $g_{i}$ with a pair of edges $q_{j},q_{k}$, it checks the current refined QIR for existing nodes with the same name. If true, it increments the weight of the edges between the nodes by one. Otherwise, it adds $q_{j}$ or $q_{k}$ or both as new nodes in the refined QIR and connects them with a weight of one.
Analysis Passes. We implement passes that analyze the QIR to identify optimization opportunities, such as gate dependencies (DependencyGraphPass), hotspot nodes (HotspotNodePass), and graph isomorphism (IsIsomorphicPass). We also implement properties passes that traverse the QIR to generate the Qernel static properties (§ 5.2) and comprise the BasicAnalysisPass and SupermarqFeaturesPass. Specifically, the former generates the key circuit properties while the latter computes the six feature vectors defined in (Tomesh et al., 2022), as explained in § 5.2. We show how the optimizer uses the information obtained from these passes in § 6.
6. Compiler Middle-end: Optimizer
QPUs comprise up to only a few 100s qubits, which sets the physical limit for circuit size and are noisy, which sets a practical limit to high fidelity execution (§ 3.1). To increase the scalability of circuits that run with high fidelity, we need a modular, extensible, and composable optimizer. Modular to support adding/removing optimization passes or changing their relative order, extensible to add new passes, and composable to chain the optimization improvements of the individual passes.
Such optimization passes are the circuit compaction techniques that, as the name suggests, reduce the circuit size, i.e., the number of qubits, rendering it executable on small QPUs and at the same time, also simplify the circuit structure, i.e., remove noisy gates. Notably, we can compose more than one of these techniques to achieve even better results. The compiler’s middle-end, the optimizer, is a composable pipeline of transformation passes that compact the QIR to increase the scalability of quantum circuits running with high execution fidelity.
@@ 6. Compiler Middle-end: Optimizer
Challenges. However, it is not trivial to implement such a pipeline. Currently, there is a plethora of individual compaction techniques that require their own sub-systems to operate, with no common infrastructure to compose them. Additionally, as we will show later, some techniques spawn an exponential number of sub-circuits (Table 1) and, after execution, require post-processing using classical hardware. We pose two questions: (1) How are the (exponentially) spawned circuits from different techniques handled? (2) How can we manage the tradeoff between fidelity improvement and exponential overheads from different techniques?
Our Approach. To this end, we design our optimizer with the goals of providing (1) a unified infrastructure for pluggable compaction mechanisms and (2) tunable knobs for configuring the tradeoff between overheads and performance improvement. For (1), we build and compose vastly different compaction techniques on the Qernel abstraction, specifically on the QIR and its refined form. For (2), we provide users with two knobs: the optimization budget (equivalent to optimization level) $b\in\mathbb{N}$ and the size to reach $s\in\mathbb{N}$, which denotes the desired post-optimization QIR size (number of qubits). Since all overheads are exponential and the exponent’s base does not make a practical difference, the single budget knob, $b$, suffices.
QIR Compaction Techniques. There are two main QIR compaction techniques: circuit divide-and-conquer and qubit reuse. In the former category, the (large) QIR is cut into smaller fragments that are executed on small QPUs, and the execution results are merged back to a single value. Circuit cutting and knitting (Peng et al., 2020; Mitarai and Fujii, 2021) belongs to this category. In our optimizer, we implement a pass that automatically cuts the QIR based on qubits (WireCuttingPass) and a pass that cuts the QIR based on gates (GateCuttingPass). To restrict the exponential overheads that scale with the number of cuts, we use the budget $b$ to cut up to $b$ times. At each cut location, the pass places a virtual gate, which must be later replaced with other gates that simulate the effects of the original (pre-cut) gate (§ 7.1).
We implement another circuit divide-and-conquer technique, also with exponential overheads, namely the QubitFreezingPass, which is limited to QAOA applications only (Ayanzadeh et al., 2023). At a high level, it removes qubits with significantly more noisy gates than other qubits, i.e., hotspot qubits, along with the gates. This reduces the number of physical qubits required by an underlying (small) QPU and greatly reduces the number of noisy gates. We use the same budget $b=3$ to remove the nodes with the highest degrees to restrict the exponential overheads.
Lastly, in the qubit reuse category, we implement the QubitReusePass, which “compacts” multiple logical qubits into one to reduce the QIR’s size (Hua et al., 2023; DeCross et al., 2022; Sadeghi et al., 2022). This process, however, increases the QIR’s depth; therefore, the tradeoff, in this case, is between QIR size and depth. To restrict the depth increase, we use qubit reuse as a last resort to achieve the user’s size requirement ($s$) or to render the Qernel executable by at least one QPU in the system.

### v2

@@ 3.1 Fidelity
Results. Our results are highlighted in Figure 3 (a). The x axis shows the circuit size as the number of qubits while the y axis shows the fidelity, where higher is better. The experiment is run on the IBM Kolkata 27-qubit QPU. For each increase in qubits, the average fidelity decreases, up to $98.9\%$ from 4 to 24 qubits. Moreover, it is physically impossible to run circuits with a size larger than 27 qubits, since we cannot map them.
Implication. NISQ devices are limited due to size and noise and, therefore, cannot be practically used for large quantum circuits, either logically because the circuit doesn’t fit in the device or the execution results would be degraded from noise-induced errors, which translates to low fidelity.
3.2 Spatial and Temporal Heterogeneity
In the classical domain, two identical CPUs perform similarly for all applications at each point in time. In contrast, QPUs exhibit differences in the layout and connectivity of qubits [26] and variations in noise errors even for QPUs of the same model, which leads to spatial performance variance. Moreover, QPUs are calibrated regularly (§ 2.2), and after each calibration, the noise properties change [61]. As a result, the execution fidelity can vary across different calibration cycles, leading to temporal performance variance.
Results. Figure 3 (b) shows a 12-qubit GHZ circuit’s fidelity on different IBM QPUs. Fidelity varies across the QPUs, with a maximum difference of 38% from best to worst. Notably, all six QPUs are of the same model (Falcon r5.11).
Figure 4 (a) shows a 6-qubit GHZ circuit’s fidelity over 120 calibration days executed on the IBM Perth 7-qubit QPU, where each data point represents a single day’s fidelity. The largest single-day difference in fidelity is 96.5%, and there are 20 instances of a single-day fidelity drop of more than 5%. Note that there is no way to predict a QPU’s future calibration data to expect such performance differences.
@@ 3.2 Spatial and Temporal Heterogeneity
3.3 Utilization
@@
Results. Figure 4 (b) shows the maximum utilization of the IBM Kolkata 27-qubit QPU for nine benchmarks while targeting at least 0.75 fidelity. No benchmark exceeds 30% utilization, while the average is 26.3%. Higher fidelity targets would yield even lower utilization and vice-versa.
Implications. There is a tradeoff between QPU utilization and fidelity. In general, the lower utilization, the higher fidelity, and vice-versa. In contrast to the classical domain, the tension between these metrics is vastly larger.
3.4 QPU Load Imbalance
The quantum cloud faces QPU load imbalance. The root cause is spatiotemporal heterogeneity (§ 3.2), combined with the manual QPU selection offered by the current quantum cloud model [32]. These naturally motivate users to select the highest fidelity QPU(s) for that calibration cycle [61].
@@ 3.4 QPU Load Imbalance
Implications. Load imbalance leads to long waiting times for the users and thus, low quality of service. However, the performance difference does not always justify the load differences between QPUs. For instance, the 12-qubit GHZ circuit in Figure 3 (b) performs 1.1$\times$ better on IBM Hanoi than IBM Cairo, yet the former exhibits 57$\times$ higher load.
4 Overview
We propose QOS, a cloud operating system for quantum computing. In this context, we define a quantum cloud operating system as a systems software layer that transparently manages quantum jobs and resources efficiently w.r.t. job fidelity and waiting times (user’s goals) and QPU utilization and load-balance (cloud operator’s goals). Therefore, QOS strives for three design goals: (1) QOS’s architecture should be general and modular, cleanly separating mechanism from policy while enabling both cross-layer coordination and within-layer optimizations. (2) QOS should enable the execution of large quantum jobs with high fidelity. (3) QOS should be resource efficient by achieving high QPU utilization and balancing QPU load to minimize queue waiting times.
Figure 5: QOS overview (§ 4): QOS consists of four main components: the error mitigator, estimator, multi-programmer, and scheduler.
4.1 The QOS Architecture
@@
Figure 5 shows the overview of our system’s design. QOS comprises a layered architecture that consists of an API and four main components: the error mitigator, the estimator, the multi-programmer, and the scheduler, which we detail next.
@@
Qernel abstraction. QOS implements a wide range of mechanisms with different abstraction requirements, from error mitigation to scheduling level. To enable the composability of these mechanisms in a unified architecture, we propose the Qernel abstraction that acts as a common denominator for the QOS mechanisms to apply their policies.
QOS API. The QOS API abstracts away the underlying complexity of the noisy and heterogeneous quantum resources by exposing hardware-agnostic functions for configurable quantum job execution. The API transforms the user’s input circuits into Qernels for QOS to apply its mechanisms.
Error mitigator. To increase the fidelity of quantum programs, the error mitigator applies pre- and post-processing error mitigation techniques. However, there exists an abundance of such techniques, and they typically incur high runtime overheads that scale with the amount of error mitigation used. Thus, the error mitigator automatically selects and composes a subset of techniques given budget constraints to limit the associated runtime costs.
Estimator. To make informed scheduling decisions in the landscape of QPU spatiotemporal performance variance, we require accurate fidelity estimations that do not depend on expensive simulations or trial runs. The estimator leverages analytical models to predict the execution fidelity on the underlying QPUs in a scalable manner.
Multi-programmer. There is an inherent tradeoff between QPU utilization (proxy of program size) and fidelity. To increase QPU utilization without significantly compromising fidelity, the multi-programmer spatially multiplexes multiple quantum programs on a single QPU with careful consideration for fidelity performance loss.
Scheduler. There is a fundamental tradeoff between job waiting times and fidelity that arises from the underlying spatial performance variance. The scheduler trades minimal fidelity for significant waiting time improvement by optimizing for this conflicting objective.
4.2 QOS Programming Model
We introduce the QOS programming model designed to abstract away the underlying complexity of managing heterogeneous and noisy quantum resources. QOS exposes hardware-agnostic APIs and leverages the Qernel unified abstraction that acts as a common denominator across its components to enable the application of its diverse mechanisms.
QOS APIs. Table 1 shows the QOS APIs that abstract away quantum job execution and resource management. To run a quantum circuit, users simply call run with the circuit and configuration options, such as the error mitigation budget, which controls the runtime overheads and thus cost (in \$). This function returns a unique job id, $jID$. Users can call result and status with the returned $jID$ to retrieve the execution results and status, respectively. Last, users call backends and backend_props to retrieve the available QPUs and their properties.
Table 1: QOS programming API.
QOS API Description
run(circ, cnfgs) Run circuit with config. options.
results(jID) Retrieve the job results.
status(jID) Retrieve the job status.
backends() Retrieve the available backends.
backend_props(bID) Retrieve the backend properties.
The Qernel abstraction. The Qernel implements data structures that store the static and dynamic properties of the quantum job. Specifically, to apply error mitigation techniques optimally, we must leverage the characteristics of the quantum circuit, such as the circuit’s size (number of qubits), depth, number and types of 2-qubit gates, the number of measurements, etc. Additionally, we include the six Supermarq features vectors [76] since they are potentially useful for heuristic-based optimizations or regression-based prediction models [59]. This information is described as the static properties of the Qernel. The dynamic properties include the Qernel’s execution status (done, failed, running, scheduled), the estimator’s output, i.e., fidelity estimations (§ 6), and the final post-processed results.
@@ 4.2 QOS Programming Model
System workflow. Users submit a large circuit using the QOS API 1. Then, QOS transforms the circuit into a Qernel, the common denominator of QOS, and submits it to the error mitigator 2. This component outputs error-mitigated Qernels and submits them to the estimator 3. The estimator predicts the fidelity of running the Qernel(s) on the QPUs to guide scheduling 4. Next, the multi-programmer bundles Qernels with low utilization and sends them to the scheduler 6, who schedules and runs the bundled Qernels, optimizing for maximal fidelity and minimal waiting times 6. After the execution, the bundled results are unbundled by the multi-programmer into separate results and are sent to the error mitigator for post-processing 8. Finally, the post-processor synthesizes the final results and returns them to the user 9.
@@
Figure 6: Error mitigator pre-processor workflow (§ 5.1). The initial Qernel has 7 qubits and 12 gates, and budget $b=3$. (a) We identify $q_{3}$ as a hotspot node with a degree of 6. (b) We remove this node, which reduces the qubit and gate counts to 6. (c) We remove two gates, giving two fragments of 3 qubits each. (d) The budget is depleted, so we use qubit reuse to further reduce the qubits by two.
5 Error Mitigator
Quantum computers are characterized by hardware and environmental noise errors, which hinder their practicality and scalability (§ 3). The error mitigator applies pre- and post-processing to the Qernels and execution results, respectively, to mitigate these noise errors.
@@ 5 Error Mitigator
Challenges. Currently, there is a plethora of individual error mitigation techniques that require their own sub-systems to operate, with no common infrastructure to compose them. However, composability is essential for fidelity improvement since error mitigation techniques can complement each other, if stacked together in the correct order, to further improve performance [11, 9]. Additionally, some techniques spawn an exponential number of sub-circuits and, after execution, require post-processing using classical resources [55, 44]. There are two main challenges in efficiently leveraging error mitigation: (1) Which error mitigation techniques must be used and in which order? (2) How to balance the tradeoff between fidelity improvement and exponential overheads w.r.t. runtime and resources?
Key ideas. The error mitigator automatically applies error mitigation techniques, abstracting this complexity away from the user. The key idea is that not all gates and qubits are equal; some of them are noise hotspots, lowering the fidelity significantly more than others. We constrain the exponential overheads by (1) setting an error mitigation budget $b$ to be spent, and (2) greedily applying error mitigation techniques on the hotspots in the order that maximizes fidelity improvement the most with the least added overheads.
5.1 Pre-Processor
The pre-processor first analyzes the Qernel to generate important metadata, identifies the error mitigation opportunities, and then prepares the Qernel for the respective techniques.
Qernel analysis. Qernel analysis generates two main pieces of information that guide the application of error mitigation techniques: (1) Qernel metadata and (2) optimization opportunities such as hotspot qubits or gates, i.e., qubits and gates that can be removed to reduce noise errors significantly. Specifically, we generate the key Qernel static properties and the six SupermarQ feature vectors as stated in § 4.2.
Error mitigation techniques. We focus on two existing error mitigation techniques: circuit divide-and-conquer and qubit reuse, since these techniques increase fidelity and enable the execution of large quantum jobs on small(er) QPUs. However, QOS supports any technique that achieves the same goals [75, 10, 43, 15].
At a high level, techniques in the former category divide (large) circuits into smaller fragments that can be executed on small QPUs, and the fragmented execution results are post-processed back to a single value. The advantages are two-fold: (1) We can scale the size of circuits that are executable on the current small QPUs beyond the H/W qubit limit, and (2) we can improve fidelity since smaller and less complex circuits achieve higher fidelity (§ 2.2, § 3.1). Unfortunately, these techniques incur exponential quantum and/or classical overheads with respect to the number of divisions; therefore, they must be used conservatively.
More specifically, circuit divide-and-conquer comprises circuit cutting and knitting [55, 44] and qubit freezing [4]. Circuit cutting can simplify the circuit’s structure by virtualizing noisy non-local gates into (less noisy) local gates [44] or cutting a qubit (wire) in the temporal dimension to shorten its runtime [55]. Thus, circuit cutting can improve the circuit’s width, depth, and number of non-local gates at once (§ 2.2). During knitting, the original (full) circuit results are reconstructed through classical post-processing of the subcircuit results. We refer interested readers to CutQC [73] and QVM [78] for details on how circuit cutting works in practice. Comparably, qubit freezing identifies qubits with significantly more connections to other qubits (hotspots) and partitions the circuit by replacing the hotspot qubits with binary values, effectively dropping their noisy non-local gates. Since every qubit can be replaced by a binary value, this process generates $2^{m}$ smaller circuits for $m$ frozen qubits.
In both cases, we automatically find the cut locations that achieve the smallest sub-circuits possible with the fewest cuts possible. To maximize the synergy between techniques, we first apply qubit freezing to remove multiple non-local gates at once with a relatively small cost, and then we greedily apply circuit cutting. To restrict the exponential overheads that scale exponentially with the number of cuts, we use the budget $b$ to cut up to $b$ times. Typically, we set the budget $b=3$ by default, and the overheads scale as $O(2^{b})-O(8^{b})$, depending on the divide-and-conquer technique.
On the other hand, qubit reuse (also referred to as recycling) reduces qubit requirements by reusing physical qubits after resetting them [29, 18, 34]. Once a qubit’s role in computation is complete, it is measured, reset to a known state (typically |$\ket{0}$), and reassigned to another logical qubit later in the circuit. This approach enables more efficient use of limited hardware resources and allows execution of larger circuits than the QPU’s qubit limit. This method, however, increases the circuit depth (i.e., execution duration), which can lead to quantum decoherence errors (§ 2.2); therefore, the tradeoff, in this case, is between circuit size and runtime. To restrict the runtime increase, we use qubit reuse as a last resort to render the Qernel executable by at least one QPU in the system (i.e., the optimized Qernel fits at least one QPU’s size constraints).